OpenStudy (anonymous):

How would you simplify (5a+4)(5a-b)

OpenStudy (anonymous):

you can't simplify it you can multiply it though

OpenStudy (anonymous):

\[(5a+4)(5a-b)=5a\times 5a-5a\times b+4\times 5a-4\times b\] is a start

OpenStudy (anonymous):

Thank You! Also would (-3/2x^4)(y^0) the answer to this be -3x^4-y?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\(y^0=1\)

OpenStudy (anonymous):

Would the answer be 45a-5ab-4b?